# Exploring limits of dipolar quantum simulators with ultracold molecules

### v1

@@ Abstract
We provide a quantitative blueprint for realizing two-dimensional quantum simulators employing ultracold dipolar molecules or magnetic atoms by studying their accuracy in predicting ground-state properties of lattice models with long-range interactions. For experimentally relevant ranges of potential depths, interaction strengths, particle fillings, and geometric configurations, we map out the agreement between the state prepared in the quantum simulator and the target lattice state. We do so by separately calculating numerically exact many-body wave functions in the continuum and single- or multi-band lattice representations, and building their many-body state overlaps. While the agreement between quantum simulator and single-band models is good for deep optical lattices with weaker interactions and low particle densities, the higher band population rapidly increases for shallow lattices, stronger interactions, and in particular above half filling. This induces drastic changes to the properties of the simulated ground state, potentially leading to false predictions. Furthermore, we show that the interplay between commensurability and interactions can lead to quasidegeneracies, rendering a faithful ground state preparation even more challenging.
Introduction — The realm of ultracold atomic and molecular systems has become a pillar for investigating fundamental quantum physics and simulating exotic many-body systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. Recent advancements in controlling and manipulating magnetic atoms [12, 13, 14, 15] and dipolar molecules [16, 17, 18, 19, 20, 21, 22] have considerably extended the already very diverse array of realizable quantum systems to those that operate under long-range interactions, with remarkable achievements such as creating molecular dipolar Bose-Einstein condensates [23], dipolar supersolid phases [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36], and dipolar quantum magnets [37, 38, 39]. Moreover, when coupled to optical lattices, ultracold dipolar particles can simulate long-ranged lattice Hamiltonians that mimic those encountered in condensed matter or high-energy physics that host a rich and unconventional phenomenology not observed in traditional short-range systems, but defy analytical treatment.
The tremendous technical possibilities offered by ultracold dipolar quantum simulators, though, raise the question of their accuracy. In fact, while the qualitative quantum simulation of lowest-band lattice models has been amply showcased, its quantitative correctness can be strongly affected by interactions, in particular long-ranged ones [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50]. It is of paramount importance to clarify this issue to capitalize on the rapid technical advancements offered by the next generation of quantum simulators and reap the rich field of long-range interacting systems. As the Bose-Hubbard model [51, 52, 53] (BHM) has been at the forefront of the quantum simulation revolution due to its fundamental applications in strongly correlated physics [54, 55, 56, 57, 58, 59, 60, 61, 62, 63], in this work we focus on the quantum simulation of its dipolar counterpart [64, 65, 66, 67, 68] that has lead to many no less groundbreaking discoveries [69, 70, 71, 72, 73, 74, 75]. We provide a systematic and quantitative comparison between the physics realized in the continuum experiment and the one observed in the corresponding lattice model. To quantify the agreement between the two descriptions, we separately calculate ground-state energies, density distributions and fidelities between many-body wave functions. This allows us to precisely delineate the experimentally relevant ranges of potential depths, interaction strengths, particle fillings, and geometric configurations where quantum simulation is accurate. Our results show that the accuracy rapidly deteriorates for shallow lattices, stronger interactions, and filling fractions above half filling. The discrepancy can be particularly high for incommensurate geometries with quasidegenerate ground states and can even lead to drastically different predictions for the density distribution. Our study provides a comprehensive chart to assess the accuracy of dipolar quantum simulators and should act as a guideline for experimental realizations. More crucially, it exemplifies the potential of such systems to realize more complex multiband lattice models hosting exotic physics.
@@
Figure 1: Accuracy of 2D dipolar quantum simulators in different parameter regimes. (a)-(i) Many-body fidelity as a measure for the accuracy of a single-band dipolar Bose-Hubbard model quantum simulator, for varying lattice depth $A_{0}$, interaction strength $V$ and filling fraction $\nu$. The labels of (b) apply to all panels.
@@
where $L_{0}$ is the distance between two neighboring minima (unit of length) and $E_{R}$ is the recoil energy of the optical lattice (unit of energy). The optical lattice has hard-wall boundaries that restrict the number of minima to exactly $S_{x}$ in $x$ direction and $S_{y}$ in $y$ direction ($S_{i}\in\mathbb{N}$), thereby recreating a lattice with $S_{x}\times S_{y}$ sites  [^1] For an even number of sites in the $j$ direction, we need to set $\phi_{j}=\pi/2$. For the dipole-dipole interactions (DDI), we assume a strong transverse harmonic confinement which suppresses the divergence [77, 78, 79, 80] and yields the regularized form
@@
If the quantum simulator is used to study a single-band lattice model, qualitative discrepancies can arise when the population of higher bands is non-negligible. To discern this, we consider dipolar BHM in the lowest-band approximation (1BDBH) and the “one-and-a-half band” approximation (1.5BDBH), where one band is used in one spatial direction and two bands are employed in the other. To obtain the ground state of the lattice Hamiltonian we perform exact diagonalization with the QuSpin Python library [97, 98, 99]. We then use the Wannier function basis to reconstruct continuum versions |$\ket{\Psi_{\mathrm{1/1.5BDBH}}}$ with corresponding density $\rho_{\mathrm{1/1.5BDBH}}(\mathbf{x})$. This procedure in turn allows us to compute many-body fidelities ${}_{\mathrm{C}}\innerproduct{\Psi}{\Psi}_{\mathrm{1/1.5BDBH}}$ to quantify the agreement between continuum and lattice descriptions [81].
Results — We begin by considering systems described by $M=S$ orbitals and Wannier functions (lowest-band), respectively. Fig. 1 offers an overview of the fidelity across different values of optical lattice depth, DDI strength, and filling. Each panel has been obtained from a specific lattice size and geometry ($3\times 3$, $3\times 4$, $4\times 4$ or $5\times 5$) and particle number ($N=3$ to $N=6$[81]. This variety should allow us to make general statements about the nature of the continuum-lattice mapping. The overall trend of Fig. 1 is that the fidelity decreases for shallower lattices (lower $A_{0}$) and stronger DDIs (higher $V$). The first observation is in accordance with general expectations that the tight-binding mapping is less precise for shallower lattices, since the interband gap is generally proportional to $A_{0}$. The second observation is more intriguing and is due to the long-range nature of the DDIs. As $V$ is increased, so does the interaction energy and more and more coupling terms between distant sites have to be accounted for in the lattice picture. These not only include single-band density-density interactions, but also density-assisted tunneling terms across sites and bands, which lead to a larger projection of the continuum ground state onto excited lattice bands. As a consequence, the fidelity of the continuum state with the single-band lattice state is reduced. This finding is compatible with previous observations in 1D geometries [50].
The increasing discrepancy between continuum and lattice calculations in the low $A_{0}$ / high $V$ regimes can also be evinced by comparing the energy of the corresponding ground states, shown in Fig. 2. Firstly, we find that the continuum ground state energy is systematically lower than the corresponding lattice one. This is expected since the continuum description is more fundamental than the mapping to the lattice. Moreover, as for the fidelity, the difference in energy increases with stronger interactions and shallower lattices, suggesting a less compatible description in these two cases.
@@
While $A_{0}$ and $V$ play an important role in setting the accuracy of the quantum simulator, from both Fig. 1 and Fig. 2 we see that the filling fraction is the crucial quantity that mostly disrupts it. For lower filling fractions up to 0.25, the agreement is excellent with a fidelity well above 99% for most of the phase diagram. For slightly larger filling fractions up to 0.44, the fidelity remains very high in most regions and only drops to around 90% at low lattice depths ($A_{0}\lessapprox 6E_{r}$) and/or very strong interactions ($V\gtrapprox 7E_{R}$). However, filling fractions above 0.5 see a rapid decline in the fidelity to values below 50%. Interestingly, this can also occur for deep lattices with $A=10E_{R}$. The reason for this decrease should again be attributed to a sharp increase in the interaction energy – this time by forcing more particles into the same space – which grants particles the ability to overcome the band gap and excite states in higher bands, thereby reducing the lowest-band population. The dipolar quantum simulator at high densities is thus definitely breaking out of the single-band picture.
@@
To quantify the contributions from higher bands, in Fig. 3 we present more detailed fidelity calculations where the 1BDBH and the 1.5BDBH are directly compared with each other for the $3\times 3$ geometry. In 1BDBH calculations, the fidelity drops monotonically for decreasing $A_{0}$ and increasing $V$ as previously seen. This worsening is particularly evident for higher filling states with $N=5$ and $N=6$, where the fidelity reaches low points of 51% and 14% respectively. However, when more bands are taken into account, the fidelity improves across all fillings. The improvement is particularly striking for $N=6$ ($\nu=0.67$). This is because the single-band lattice model predicts a drastically different ground state (with a checkerboard pattern) than the one obtained from the 1.5-band model (with a striped pattern), as can be seen in the insets of Fig. 3(d). This finding delivers two important messages. On the one hand, if the target simulation model is restricted to the single-band model, the quantum simulator can dramatically fail to obtain the correct ground states at high fillings and thus lead to wrong physical predictions. On the other hand, the physical setup is already capable of generating richer physics of more complex higher-band models without additional fine tuning.
@@
Another key factor in the accuracy of the quantum simulator, less evident from the aggregated filling data, is system geometry. Depending on the interplay between particle number, repulsion strength, and lattice geometry, incommensurability might occur, i.e. non-integer filling for each occupied site. This can lead to quasidegenerate ground states due to spontaneous lattice symmetry breaking and thus make populating the true ground state an arduous task in an experimental realization. An overview of the quasidegeneracies emerging in the lattice system is presented in Fig. 4, where we plot the gap between the ground state and the first excited state for all the geometries considered at the four extremal values in our $(V,A_{0})$-parameter range. We can immediately distinguish two clusters of states. The clustering itself is essentially independent of the values of $V$ and $A_{0}$, thereby revealing that it is rather determined by the underlying geometry. However, within each cluster $V$ has opposite effects (stabilizing vs. destabilizing). The first cluster consists of $N=4,5$ in the $3\times 3$ and $N=4$ in the $4\times 4$ lattice, i.e. geometries that can accommodate the ground state in the most symmetric configuration in the localized limit (four isolated particles in the corners and none or one in the center, respectively). These states preserve all discrete symmetries of the lattices, such as $C_{4}$ rotations, reflections, and inversions, which makes them very stable as confirmed by a large gap to excited states. The second cluster, instead, consists of geometries with incommensurability. Examples include all the $N=3$ states and the $N=5$ state in the $4\times 4$ lattice. For these states, the localizing forces (lattice potential and repulsions) clash with the available minima compatible with the lattice symmetries. In milder cases (e.g. $N=3$, $5\times 5$ lattice), this favors a more delocalized ground state which does accommodate the symmetries but can be easily excited to similar states of quasidegenerate energies  [^2] The half-filled $S_{x}\times S_{y}=3\times 4$ case despite being commensurate was added to the incommensurate cluster due to its completely degenerate ground state.. In more severe cases (e.g. $N=6$, $3\times 3$ and $3\times 4$ lattices), a total breaking of rotational symmetries occurs, leading to a ground state partitioned into multiply-degenerate sectors, e.g. two opposite checkerboard patterns, a feature known in extended BH models for states above half filling [101, 102, 103]. Thus, we reveal another pitfall of blindly employing dipolar quantum simulators, namely the danger of preparing highly degenerate states with very different density distributions and physical properties.
@@
We have presented a comprehensive analysis of the accuracy of 2D dipolar quantum simulators by performing quantitative comparisons between continuum models of experimental setups and effective lattice models in a wide range of relevant lattice depths, interaction strengths, and filling fractions. Our study highlights that ultracold dipolar systems can be a very powerful instrument to reach regimes of multiband occupation and thereby allow to probe more realistic, complex long-range interacting lattice models. Yet, the population of higher bands can drastically alter ground-state properties, leading to erroneous predictions if the target lattice model is within a single-band picture. This is exemplified for $N=6$ particles in a $3\times 3$ geometry, where the 1.5BDBH model predicts a striped state, but the 1BDBH model gives rise to a checkerboard pattern. In general, regimes of high interaction strengths, shallow lattices, and high fillings should generate the richest multiband physics. While highly symmetric states could be simulated with greater accuracy, incommensurability effects emerging from the competition between spatial geometry and localizing interactions and potentials can drastically alter the spectral landscape. As a consequence, quasidegenerate ground states can arise, which can be completely modified by the presence of higher-band contributions. Our findings highlight a fundamental challenge in faithfully preparing and controlling single-band, long-range interacting states in ultracold quantum simulators – particularly above half-filling. This should even prompt the question whether single-band descriptions are sufficient to encompass the most exotic physics of systems (e.g. in hard condensed matter) which interact via even longer ranged couplings such as Coulomb forces. On the other hand, our results showcase the enormous possibilities offered by dipolar quantum simulators to reproduce novel physics in effective multiband lattice models.
Our study, based on a rigorous, quantitative comparison between lattice and continuum descriptions, should provide a reliable blueprint for the correct realization of long-ranged quantum simulators. While we have focused on repulsive interactions in two dimensions, we expect similar accuracy problems to arise for attractive regimes and different spatial geometries, too. This is particularly important in light of discoveries of interesting new phases in attractive dipolar BHM models such as exotic quantum liquids [75, 74, 108] and self-bound quasicrystalline order [109]. In particular, a key question to address will be the treatment of singularities arising for DDIs at zero distance, which can occur in the continuum but are truncated by the lattice formulation. Our approach of quantitatively comparing continuum and lattice descriptions by calculating many-body overlaps should shed light on those issues.

### v2

# Realizing multiband states with ultracold dipolar quantum simulators

@@ Abstract
The manipulation of dipolar interactions within ultracold molecular ensembles represents a pivotal advancement in experimental physics, aiming at the emulation of quantum phenomena unattainable through mere contact interactions. Our study uncovers regimes of multiband occupation which allow to probe more realistic, complex long-range interacting lattice models with ultracold dipolar simulators. By mapping out experimentally relevant ranges of potential depths, interaction strengths, particle fillings, and geometric configurations, we calculate the agreement between the state prepared in the quantum simulator and a target lattice state. We do so by separately calculating numerically exact many-body wave functions in the continuum and single- or multiband lattice representations, and building their many-body state overlaps. Our findings reveal that for shallow lattices and stronger interactions above half filling, multiband population increases, resulting in fundamentally different ground states than the ones observed in simple lowest-band descriptions, e.g. striped vs checkerboard states. A wide range of probed parameter regimes in its turn provides a systematic and quantitative blueprint for realizing multiband states with two-dimensional quantum simulators employing ultracold dipolar molecules.
Introduction — Ultracold atomic and molecular systems have emerged as fundamental tools for exploring quantum physics and simulating complex many-body phenomena [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. The ability to control and manipulate magnetic atoms [12, 13, 14, 15] and dipolar molecules [16, 17, 18, 19, 20, 21, 22] has dramatically broadened the spectrum of quantum systems we can study, particularly those governed by long-range interactions. This has led to significant breakthroughs, including the creation of molecular dipolar Bose-Einstein condensates [23], supersolid phases [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36], and quantum magnets [37, 38, 39]. Additionally, when integrated with optical lattices, these systems allow for the simulation of long-range Hamiltonians akin to those in condensed matter and high-energy physics, unveiling unique behaviors not seen in traditional systems with short-range interactions. This progress in experimental capabilities is matched by theoretical advancements, enhancing our understanding of quantum systems’ properties and dynamics and uncovering novel phenomena such as new quantum and spin liquid phases, altered phase transition behaviors, and modified dynamics and entanglement spread [40, 41, 42, 43].
A particularly exciting development in this field is employing interactions to more readily construct complex extensions of prototypical lattice models such as the Bose-Hubbard model (BHM) [44, 45, 46], a cornerstone in the study of strongly correlated physics [47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67]. While research has largely focused on the BHM’s adaptation to long-range interactions [68, 69, 70, 71, 72] and the new quantum phases that emerge from them [73, 74, 75, 76, 77, 78], the potential of dipolar simulators to explore multiband models has received less attention. In our study, we precisely identify the conditions — spanning potential depths, interaction strengths, particle fillings, and geometric configurations — under which these multiband models can be realized. We provide a detailed comparison of the physics observed in experimental setups to that in the corresponding lattice models, assessing their congruence through ground-state energies, density distributions, and many-body wave function fidelities. Our findings reveal that for shallow lattices and stronger interactions above half filling, higher bands become rapidly populated, resulting in fundamentally different ground states than the ones observed in simple lowest-band descriptions, e.g. striped vs checkerboard states. This discrepancy is especially pronounced in incommensurate geometries with nearly degenerate ground states. Our work highlights the potential to simulate more intricate multiband lattice models with ultracold dipolar systems, offering insights for future experimental designs and a roadmap for evaluating dipolar quantum simulators’ accuracy.
@@
Figure 1: Fidelity of 2D dipolar quantum simulators and 1BDBH model in different parameter regimes. (a)-(i) Many-body fidelity as a measure for the transition of a single- to multi-band dipolar Bose-Hubbard model quantum simulator, for varying lattice depth $A_{0}$, interaction strength $V$ and filling fraction $\nu$. The labels of (b) apply to all panels.
@@
where $L_{0}$ is the distance between two neighboring minima (unit of length) and $E_{R}$ is the recoil energy of the optical lattice (unit of energy). The optical lattice has hard-wall boundaries that restrict the number of minima to exactly $S_{x}$ in $x$ direction and $S_{y}$ in $y$ direction ($S_{i}\in\mathbb{N}$), thereby recreating a lattice with $S_{x}\times S_{y}$ sites  [^1] For an even number of sites in the $j$ direction, we need to set $\phi_{j}=\pi/2$. For the dipole-dipole interactions (DDI), we assume a strong transverse harmonic confinement which suppresses the divergence [80, 81, 82, 83, 84] and yields the regularized form
@@
Long-range interactions in closed geometries can quickly lead to higher-band populations. To discern this, we consider dipolar BHM in the lowest-band approximation (1BDBH) and the “one-and-a-half band” approximation (1.5BDBH), where one band is used in one spatial direction and two bands are employed in the other. To obtain the ground state of the lattice Hamiltonian we perform exact diagonalization with the QuSpin Python library [101, 102, 103]. We then use the Wannier function basis to reconstruct continuum versions |$\ket{\Psi_{\mathrm{1/1.5BDBH}}}$ with corresponding density $\rho_{\mathrm{1/1.5BDBH}}(\mathbf{x})$. This procedure in turn allows us to compute many-body fidelities ${}_{\mathrm{C}}\innerproduct{\Psi}{\Psi}_{\mathrm{1/1.5BDBH}}$ to quantify the agreement between continuum and lattice descriptions [85].
Results — We begin by considering systems described by $M=S$ orbitals and Wannier functions (lowest-band), respectively. Figure 1 delineates a comprehensive portrayal of fidelity across a spectrum of variables including optical lattice depth, DDI intensity, and filling fractions. Each graphical representation is derived from a distinct combination of lattice dimensions and configurations ($3\times 3$, $3\times 4$, $4\times 4$, or $5\times 5$) in conjunction with a range of particle numbers ($N=3$ to $N=6$[85]. Such diversity facilitates the formulation of overarching conclusions concerning the continuum-lattice mapping paradigm. Predominantly, Fig. 1 exhibits a tendency where fidelity diminishes for shallower lattices (lower $A_{0}$) and intensified DDIs (higher $V$).
The first observation is in accordance with general expectations that the tight-binding mapping is less precise for shallower lattices, since the interband gap is generally proportional to $A_{0}$. The second observation unveils a fascinating aspect linked to the extended reach of DDIs. As $V$ is increased, so does the interaction energy and more and more coupling terms between distant sites have to be accounted for in the lattice picture. This encompasses not merely single-band density-density interactions, but also density-assisted tunneling terms across sites and bands, culminating in a larger projection of the continuum ground state onto excited lattice bands. Consequently, the fidelity of the continuum state with the single-band lattice state is reduced, a phenomenon consistent with prior studies in 1D setups [67].
The increasing discrepancy between continuum and lattice calculations in the low $A_{0}$ / high $V$ regimes is also corroborated by comparing the energy of the corresponding ground states, shown in Fig. 2. The continuum ground state energy being systematically lower than the corresponding lattice one reiterates that the continuum description is more fundamental than the mapping to the lattice. Moreover, as with the fidelity, the difference in energy increases with stronger interactions and shallower lattices, suggesting an increased population of higher bands.
@@
While $A_{0}$ and $V$ significantly influence the amount of higher-band contribution, what really affects this is the filling fraction $\nu$. Both Fig. 1 and Fig. 2 show that up to $\nu=0.25$, the lowest-band dominates the physics with above 99% fidelity across most scenarios. For slightly larger filling fractions up to 0.44, the fidelity remains very high in most regions and only drops to around 90% at low lattice depths ($A_{0}\lessapprox 6E_{r}$) and/or very strong interactions ($V\gtrapprox 7E_{R}$). However, filling fractions above 0.5 see a rapid decline in the fidelity to values below 50%. Interestingly, this can even occur for deep lattices with $A=10E_{R}$. The reason for this decrease should again be attributed to a sharp increase in the interaction energy – this time by cramming more particles – which grants them the ability to overcome the band gap and excite states in higher bands, thereby reducing the lowest-band population. The dipolar quantum simulator at high densities is thus definitely breaking out of the lowest-band picture.
@@
To quantify the contributions from higher bands, in Fig. 3 we present a detailed comparisons between the 1BDBH and the 1.5BDBH models for the $3\times 3$ geometry. In simpler 1BDBH calculations, as previously seen, the fidelity drops monotonically for decreasing $A_{0}$ and increasing $V$. Higher filling states with $N=5$ and $N=6$ exhibit the most striking fidelity drop, with low points of 51% and 14% respectively. However, by including additional bands, accuracy improves for all scenarios. This improvement is most notable for $N=6$ ($\nu=0.67$), where the two models predict completely different states as can be seen in the insets of Fig. 3(d): the ground state of the 1BDBH model exhibits a checkerboard pattern, while the one of the 1.5BDBH is instead in a striped configuration. This illustrates two key points: Approximating the dipolar quantum simulator with lowest-band models can lead to significant errors in high-density situations. Conversely, the setup is already equipped to handle richer, higher band models without extra adjustments.
@@
Another key factor that affects the properties of the ground state in the quantum simulator is system geometry. Depending on particle number and lattice geometry, strong repulsion may make the particle ordering incommensurate. This can lead to spontaneous lattice symmetry breaking which generates quasidegenerate ground states. A summary of the lattice quasidegeneracies is shown in Fig. 4, quantified as the gap between the ground state and the first excited state for all the geometries considered at the four extremal values in our $(V,A_{0})$-parameter range. Two clusters of states can be identified, with the clustering itself being essentially independent of the values of $V$ and $A_{0}$, thereby reflecting the underlying geometry. Within each cluster, $V$ has opposite effects (stabilizing vs. destabilizing). The first cluster consists of $N=4,5$ in the $3\times 3$ and $N=4$ in the $4\times 4$ lattice, i.e. geometries that can accommodate the ground state in the most symmetric configuration in the localized limit (four isolated particles in the corners and none or one in the center, respectively). These states preserve all discrete symmetries of the lattices, such as $C_{4}$ rotations, reflections, and inversions, which makes them very stable as confirmed by a large gap to excited states. The second cluster, instead, encompasses geometries with incommensurability. Examples include all the $N=3$ states and the $N=5$ state in the $4\times 4$ lattice. For these states, localizing forces such as lattice potential and repulsions clash with the available minima provided by the lattice symmetries. In milder cases (e.g. $N=3$, $5\times 5$ lattice), this favors a more delocalized ground state which does accommodate the symmetries but can be easily excited to similar states of quasidegenerate energies  [^2] The half-filled $S_{x}\times S_{y}=3\times 4$ case despite being commensurate was added to the incommensurate cluster due to its completely degenerate ground state.. In more severe cases (e.g. $N=6$, $3\times 3$ and $3\times 4$ lattices), a total breaking of rotational symmetries occurs, leading to a ground state partitioned into multiply-degenerate sectors, e.g. two opposite checkerboard patterns, a feature known in extended BH models for states above half filling [105, 106, 107]. These results demonstrate that DDI not only offer an experimental platform to realize multiband models, but in conjunction with different geometries can also produce ground-state (quasi)degeneracies that are important ingredients to achieve exotic quantum states i.e. with topological features.
@@
We have presented a comprehensive analysis of the accuracy of 2D dipolar quantum simulators by performing quantitative comparisons between continuum models of experimental setups and effective lattice models in a wide range of relevant lattice depths, interaction strengths, and filling fractions. Our study highlights that ultracold dipolar systems can be a very powerful instrument to reach regimes of multiband occupation and thereby allow to probe more realistic, complex long-range interacting lattice models. The population of higher bands can drastically alter ground-state properties. This is exemplified for $N=6$ particles in a $3\times 3$ geometry, where the 1.5BDBH model predicts a striped state (with anisotropic spatial order), but the 1BDBH model gives rise to a checkerboard pattern. In general, regimes of high interaction strengths, shallow lattices, and high fillings should generate the richest multiband physics. While highly symmetric states could be simulated with greater accuracy, incommensurability effects emerging from the competition between spatial geometry and localizing interactions and potentials can drastically alter the spectral landscape. As a consequence, quasidegenerate ground states can arise, which can be completely modified by the presence of higher-band contributions. These findings should even prompt the question whether lowest-band descriptions are sufficient to encompass the most exotic physics of lattice models systems, for instance in solid state settings, which interact via even longer ranged couplings such as unscreened or weakly screened Coulomb forces, e.g. in low-dimensional or thin-film systems. Our results showcase the enormous possibilities offered by ultracold dipolar quantum simulators to reproduce these multiband lattice models.
Our study, based on a rigorous, quantitative comparison between lattice and continuum descriptions, should provide a reliable blueprint for the correct realization of long-ranged quantum simulators. While we have focused on repulsive interactions in two dimensions, we expect a rich physics to arise for attractive regimes and different spatial geometries, too. This is particularly important in light of discoveries of interesting new phases in attractive dipolar BHM models such as exotic quantum liquids [43, 78, 112] and self-bound quasicrystalline order [113]. In particular, a key question to address will be the treatment of singularities arising for DDIs at zero distance, which can occur in the continuum but are truncated by the lattice formulation. Our approach of quantitatively comparing continuum and lattice descriptions by calculating many-body overlaps should shed light on those issues.